Solve the following; don't forget to check for extraneous roots:
SQRT(x^3) = SQRT(x^2)
A.{0,1}
B.{-1,0,1}
C.0
D.1
E.all real numbers
F.none of the above
*I answered c at first but it was wrong. Please explain your answer. Thanks in advance!

Answer by rothauserc(4718)

SQRT(x^3) = SQRT(x^2)
note that sqrt(x^3) = sqrt(x^2*x) = x*sqrt(x)
x*sqrt(x) = x
sqrt(x) = 1
x = 1
answer is D.